OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (6, 2) and (7, 4) a.-2 b.1/2 c.-1/2 d. 2

OpenStudy (anonymous):

do you know the slope formula point slope formula

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

what formulas do you know

OpenStudy (anonymous):

wait yes. y-y1=m(x-x1)

OpenStudy (anonymous):

technically the the whole idea is this 

OpenStudy (anonymous):

did i confuse you or ....

OpenStudy (anonymous):

no, i follow.

OpenStudy (anonymous):

is it d?

OpenStudy (anonymous):

well can you show me your work how you came to your answer

OpenStudy (anonymous):

2-4=m(6-7) -2=m(-1) (-2)(-1) 2

OpenStudy (anonymous):

you are supposed to get the slope first you are using the point slope formula

OpenStudy (mathstudent55):

You have two points, so you need the two-point equation of a line: A line that passes through points \((x_1, ~y_1)\) and \((x_2, ~y_2) \) has the equation: \(y - y_1 = \dfrac{y_2 - y_1}{x_2 - x_1} (x - x_1)\) Use \( (6, ~2) \) as \( (x_1, ~y_1) \) and \( (7, ~4) \) as \( (x_2, ~y_2) \). Plug in all 4 values into the two-point equation above and simplify.

OpenStudy (anonymous):

the slope formula is y - y1 ----- = m x- x1

OpenStudy (anonymous):



OpenStudy (anonymous):

i'm confused. i tried that all together and got two again.

OpenStudy (anonymous):

lol you might be correct :P i haven't tried it out actually lol

OpenStudy (anonymous):

looks good :P

OpenStudy (anonymous):

okay, thank you!

OpenStudy (mathstudent55):

\(y - y_1 = \dfrac{y_2 - y_1}{x_2 - x_1} (x - x_1) \) \(y - 2 = \dfrac{4 - 2}{7 - 6} (x - 6) \) \(y - 2 = \dfrac{2}{7} (x - 6) \) \(y - 2 = 2(x - 6) \) \(y - 2 = 2x - 12\) \(y = 2x - 10\)